OpenStudy (anonymous):

Any statistics people on here!? I need help with probability using the multiplication rule!!!

OpenStudy (anonymous):

All right, what's your question?

OpenStudy (anonymous):

I just cannot figure out how they got this answer.

OpenStudy (anonymous):

So you start by getting the NON defective ones in a fraction which is 669-160 = 509. so 509/669 are Not defective

OpenStudy (anonymous):

and then I dont know what to do after that because i've confused on why that isnt the answer.

OpenStudy (anonymous):

i'm*

OpenStudy (anonymous):

All right, so we have: \[ \frac{160}{669} \]CD's are defective. The probability of us choosing one that is not defective: \[ \frac{669-160}{669}=\frac{509}{669} \]And, continually, as we choose more CD's, we have less chance of picking all non-defective, in a row, our probability becomes: \[ \frac{_{509}P_8}{_{669}P_8}=\frac{509\cdot508\cdot507\cdot506\cdot505\cdot504\cdot503\cdot502}{669\cdot668\cdot667\cdot666\cdot665\cdot664\cdot663\cdot662}\approx0.1108 \]And, that should be our answer.

OpenStudy (anonymous):

but it says the answer is .112 :/

OpenStudy (anonymous):

how about we try another one? something a little bit different

OpenStudy (anonymous):

I know. You should have a bound of error, since some calculators use scientific notation to compute these large numbers and lose precision. But, yes, go for it.

OpenStudy (anonymous):

again the answer is given, but i cant figure out why

OpenStudy (anonymous):

I'm not sure what BOTH is referring to, so I dont know which two to multiply against eachother.

OpenStudy (anonymous):

So, we start with 48 possible cases involving an intoxicated driver, out of 166: \[ \frac{48}{166} \]We have a mutually exclusive even, since, if we pick our initial case to be intoxicated, both numbers are lowered by unity, so we have, the probability of picking a second intoxicated driver is: \[ \frac{47}{165} \]These are mutually exclusive events, so we have: \[ P(A\wedge B)=\frac{48}{166}\cdot\frac{47}{165}\approx0.082365 \]

OpenStudy (anonymous):

So if the question asked if 3 different deaths were selected without replacement it would be (48/166)*(47/165)*(46/164) =.023?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

I get that one! thank you. how about another one, similar situation

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

so this one has intercepting numbers, and it says WITH replacement, so the total number doesnt change. but when i tried to number it out, it's not the right answer.

OpenStudy (anonymous):

So, what is the probability of picking a Group B? \[ \frac{39}{249} \]And, now that we have our Group B, what is the probability of picking \(\text{Rh}^+\) from within that group? \[ \frac{37}{39} \]Which means, for one person: \[ \Pr(B \wedge \text{Rh}^+)=\frac{37}{39}\cdot\frac{39}{249}=\frac{37}{249}\approx 0.1485944 \]This means, since we are drawing with replacement, that drawing two people is: \[ \left(\frac{37}{249}\right)^2\approx0.02208 \]Which is what we wished to find.

OpenStudy (anonymous):

wait i see you have 37/39 * 39/249 = 37/249

OpenStudy (anonymous):

why dont you use the 37/39?

OpenStudy (anonymous):

also, how do you configure the second portion of the question?

OpenStudy (anonymous):

For the latter case, drawing without replacement, we go by the same process, but compute each probability individually, and, instead of squaring the probability, we multiply the last two in the end. As for the \(\frac{37}{39}\)... it is used. It's just when multiplied out it comes out to such.

OpenStudy (anonymous):

The second portion? Well, we know the probability of picking two people, with replacement, is the same, so we simply multiply the result with itself.

OpenStudy (anonymous):

hmm okay so i will attempt to see if this is right! thank you LolWolf!

OpenStudy (anonymous):

have time for another? lol

OpenStudy (anonymous):

I have one more

OpenStudy (anonymous):

Sure, go for it.

OpenStudy (anonymous):

need help with B and C

OpenStudy (anonymous):

For B, both events are mutually exclusive, so: \[ \Pr(A\wedge B)=\Pr(A)\cdot\Pr(B)=0.063\cdot0.063=0.003969 \]

OpenStudy (anonymous):

okay that makes sense

OpenStudy (anonymous):

For C, we have to compute the inverse (because the probability of multiplying both is actually interpreted as "what is the probability of both alarm clocks going off?) We know: \[ \Pr(\neg A)=1-\Pr(A) \]So, we have that, the probability of both *not* sounding is: \[ \Pr(A\wedge B)=0.003969 \]Therefore, the probability of at least ONE sounding is: \[ \Pr(\neg(A\wedge B))=1-\Pr(A\wedge B)=1-0.003969=0.996031 \]For these, I recommend reading De Morgan's laws.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/De_Morgan%27s_laws

OpenStudy (anonymous):

thank you!! i'll check that out right now! thank you for all your help. you have been great!!!!

OpenStudy (anonymous):

Sure thing